# Iterative Deepening Sampling for Large Language Models

### v1

@@ 7 Conclusions
In this paper, we introduce Iterative Deepening (ID) Sampling, a simple yet effective sampling algorithm designed to adaptively allocate computational budget for sequential revisions in large language models (LLMs). We demonstrate that ID-sampling effectively enhances inference-time performance by improving the pass rate on challenging mathematical reasoning tasks across various models. Our results indicate that while current models exhibit strong self-correction capabilities, they still have significant room for improvement in determining when to trigger self-correction themselves without human intervention.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.

### v2

@@ 7 Conclusions
In this paper, we introduce Iterative Deepening (ID) Sampling, a simple yet effective sampling algorithm designed to more efficiently scaling test-time compute compared to regular sampling. We demonstrate that ID-sampling effectively enhances inference-time performance by improving the pass rate on challenging mathematical reasoning tasks across various reasoning models on math problems. We also found that ID-sampling is less effective on non-reasoning models given their weaker self-correction capability. Our results indicate that while current models exhibit strong self-correction capabilities, they still have room for improvement in determining when to trigger self-correction themselves without human intervention, which might not be sufficiently solved by the current reinforcement learning paradigm.
Appendix A Additional Discussions
Potential Extension to Search-based Test-time Scaling
In this paper, our study primarily focuses on ID-sampling without integrating it into tree-search algorithms. However, we emphasize that our proposed framework can be directly extended to methods such as beam search and Monte Carlo Tree Search (MCTS). In these cases, the budget in Algorithm 1 can be directly interpreted as the computational budget in MCTS or the number of iterations in beam search. One challenge in this extension is that self-correction mechanisms reduce the reliance on early model outputs being correct, distinguishing the need for a new generation of PRMs from previously released ones. However, future researchers who develop or have access to more accurate PRMs that explicitly model the self-correction process can leverage our ID-sampling framework to enhance the reasoning capabilities of fixed models.
Choice of Datasets
In our experiments, we focus exclusively on mathematical reasoning datasets. While test-time scaling techniques can, in principle, be extended to more general domains, such extensions are not directly applicable in our setting. From a methodological standpoint, the injection of trigger sentences is not well suited to code generation tasks. In such cases, the insertion of additional tokens can introduce syntactic or semantic errors. More broadly, this work aims to develop a more efficient test-time scaling strategy. One promising direction is to leverage high-quality reward models to aggregate multiple responses, as demonstrated in our approach. However, existing open-source reward models are predominantly designed for mathematical problem solving, and robust, general-purpose reward models are currently lacking. Additionally, popular benchmarks such as MMLU (Wang et al., 2024) consist primarily of single-choice questions. For these tasks with a limited number of options, evaluation metrics like Pass@$n$ (for $n>1$) are not meaningful, as simply generating diverse final options can trivially improve performance. As a result, most standard non-mathematical datasets are not well aligned with the objectives of this study. We leave the exploration of ID-sampling on less commonly used datasets as an avenue for future research.
Appendix B Linguistic Markers
In the paper, we have presented the occurrence frequencies of the linguistic markers as our motivation. We provide the full list of such markers here, which is the same as (Wang et al., 2025): ”however”, ”wait”, ”alternatively”, ”hmm”.